# Bidirectional Human Interactive AI Framework for Social Robot Navigation

### v1

@@ Abstract
Trustworthiness is a crucial concept in the context of human-robot interaction. Cooperative robots must be transparent regarding their decision-making process, especially when operating in a human-oriented environment. This paper presents a comprehensive end-to-end framework aimed at fostering trustworthy bidirectional human-robot interaction in collaborative environments for the social navigation of mobile robots. Our method enables a mobile robot to predict the trajectory of people and adjust its route in a socially-aware manner. In case of conflict between human and robot decisions, detected through visual examination, the route is dynamically modified based on human preference while verbal communication is maintained. We present our pipeline, framework design, and preliminary experiments that form the foundation of our proposition.
@@ I Introduction
The trustworthiness of the AI algorithms is generally stressed by conducting an exhaustive search on the state space of the models [4], [5], segmenting plans for the monitoring purposes [6], [7], [8], and failure explanation [9], [10]. In these settings, a human expert is necessary for offline validation of decisions, while providing online explanations during operation for non-experts is an area that has received limited study.
@@ II-B Trustworthiness in Robotics
Figure 2: A potential scenario of social robot navigation is depicted. The robot adjusts its original path (current robot motion), taking into account the predicted trajectory (forecasted human trajectory) of an encountered human. However, the robot initiates movement toward the current position of the human (socially aware robot trajectory), which may induce fear and erode trust. In the second figure, the proposed system is illustrated. To address the discomfort caused by the robot’s unpredictable movements, the proposed system explains the trajectory, expecting guidance in return (trustworthy robot trajectory). As a result, the respondent human understands the robot’s intentions, provides guidance, and continues along his walk.
@@ IV Preliminary Experiments
The proposed system has localization, tracking, trajectory forecasting, and gesture detection components. While the entire navigation system is not yet complete, we have initiated experimentation with selected components. In the following, we present the preliminary results obtained from these components. Integration of these individuals into a full system will be addressed in subsequent stages of our research.
@@
To estimate trajectories, we applied our encoding approach to the Trajnet++ dataset. Fig. 6 shows an example of the visual output of the LSTM-based autoencoder on the test set. The LSTM encoder effectively represents the trajectories in the latent space and can reconstruct them. It is important to note that each trajectory will be encoded to the latent space and then used as the nodes of the graph that GAT will process.
As performance metrics, we will use Average Displacement Error (ADE) between each position step, Final Destination Error (FDE) considering only the last position accuracy, and Relative Mean Error (RME) to eliminate the error accumulation. As a distance metric, we will use Euclidean distance ($L_{2}$).
IV-A Real-world Data Collection and Management
The proposed system will undergo validation using custom data collected from a smart factory environment in TUBITAK Campus where mobile robots and humans work together on the same assembly line. Industry-standard Omron LD60 autonomous robots will be used for the mobile robot component. The trust between the robot and the human will be validated by both experts and non-experts. This study strictly adheres to General Data Protection Regulation (GDPR) [45] guidelines to protect the privacy and anonymity of data collected. All data will be anonymized, securely stored, and only accessible by authorized personnel. Participants will give informed consent and can withdraw at any time, ensuring ethical standards and GDPR compliance are met.
@@ V Conclusion and Future Works
To assess the AI module’s trustworthiness, we plan to survey individuals in smart factory settings regarding their comfort with the robot’s vocal communication and the effectiveness and variety of hand gestures for controlling the robot. Participants will rate their comfort and satisfaction with the robot’s vocal responses using a list of candidate sentences. Furthermore, participants will redirect the robot and offer feedback on the ease of navigating it using hand gestures in various scenarios.

### v2

@@ Abstract
Trustworthiness is a crucial concept in the context of human-robot interaction. Cooperative robots must be transparent regarding their decision-making process, especially when operating in a human-oriented environment. This paper presents a comprehensive end-to-end framework aimed at fostering trustworthy bidirectional human-robot interaction in collaborative environments for the social navigation of mobile robots. In this framework, the robot communicates verbally while the human guides with gestures. Our method enables a mobile robot to predict the trajectory of people and adjust its route in a socially-aware manner. In case of conflict between human and robot decisions, detected through visual examination, the route is dynamically modified based on human preference while verbal communication is maintained. We present our pipeline, framework design, and preliminary experiments that form the foundation of our proposition.
@@ I Introduction
While the concept of inter-human trust is quantitatively examined [4], the trustworthiness of the AI algorithms is generally stressed by conducting an exhaustive search on the state space of the models [5], [6], segmenting plans for the monitoring purposes [7], [8], [9], and failure explanation [10], [11]. In these settings, a human expert is necessary for offline validation of decisions, while providing online explanations during operation for non-experts is an area that has received limited study.
@@ II-B Trustworthiness in Robotics
Figure 2: A potential scenario of social robot navigation is depicted. The robot adjusts its original path, taking into account the predicted trajectory of an encountered human. However, the robot initiates movement toward the current position of the human, which may induce fear and erode trust. In the second figure, the proposed system is illustrated. To address the discomfort caused by the robot’s unpredictable movements, the proposed system explains the trajectory, expecting guidance in return.
@@ IV Preliminary Experiments
The proposed system has localization, tracking, trajectory forecasting, social navigation, and gesture detection components. While the entire navigation system is not yet complete, we have initiated experimentation with selected components. In the following, we present the preliminary results obtained from these components. Integration of these individuals into a full system will be addressed in subsequent stages of our research.
@@
To estimate trajectories, we applied our encoding approach to the Trajnet++ dataset. Fig. 5 shows an example of the visual output of the LSTM-based autoencoder on the test set. The LSTM encoder effectively represents the trajectories in the latent space and can reconstruct them. It is important to note that each trajectory will be encoded to the latent space and then used as the nodes of the graph that GAT will process. As performance metrics, we will use Average Displacement Error (ADE) between each position step, Final Destination Error (FDE) considering only the last position accuracy, and Relative Mean Error (RME) to eliminate the error accumulation. As a distance metric, we will use Euclidean distance ($L_{2}$). To evaluate the social navigation module of our system, we will compare it with CADRL [25] as a baseline, considering the aforementioned metrics.
@@ V Conclusion and Future Works
To assess the AI module’s trustworthiness, we plan to survey individuals in smart factory settings regarding their comfort with the robot’s vocal communication and the effectiveness and variety of hand gestures for controlling the robot. Additionally, as part of the ablation study, our system will be compared in the survey to a version where the human does not redirect the robot at all. Participants will rate their comfort and satisfaction with the robot’s vocal responses using a list of candidate sentences. Furthermore, participants will redirect the robot and offer feedback on the ease of navigating it using hand gestures in various scenarios.